OpenStudy (anonymous):

Tell whether each function is linear. If so, graph the function. 2x + 3y = 5

OpenStudy (agreene):

So, to start just solve this guy for y so you get something you know how to graph, 2x+3y=5 subtract your x's 3y=5-2x divide by 3 y= 5/3 - 2x/3 graph that. And yes, it is Linear.